OpenStudy (anonymous):

Paul’s Pool Service charges $340 to open a pool and charges $28 each week for pool maintenance. Nelson’s Pool Service charges $280 to open a pool and charges $32 each week for pool maintenance. The system that models this situation is given, where c is the cost of pool maintenance and w is the number of weeks. c = 340 + 28w c = 280 + 32w The solution to the system is (15, 760). Which interpretation correctly describes the solution to the system of equations? A. At 15 wk of pool cleanings, a customer will reach the maximum payment of $535. B.

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

@lsherron

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@WWWHHHAAATTT?

OpenStudy (wwwhhhaaattt?):

Wellllll Very confusing

OpenStudy (anonymous):

yep thats why i am here

OpenStudy (phi):

if we used x and y instead of w and c we would have these two equations y = 28x+340 y = 32x +280 those are the equations for 2 different lines. the lines will meet at some point (i.e. some (x,y) pair) that point is the solution of the system

OpenStudy (wwwhhhaaattt?):

Im not the best in math. Well I suck at math but ill try to help.

OpenStudy (phi):

In your problem they tell you the solution is (15, 760) any idea with number is "c" and which is "w" ?

OpenStudy (anonymous):

hang on the hole problem is not up there i will post it on the comments

OpenStudy (anonymous):

Paul’s Pool Service charges $340 to open a pool and charges $28 each week for pool maintenance. Nelson’s Pool Service charges $280 to open a pool and charges $32 each week for pool maintenance. The system that models this situation is given, where c is the cost of pool maintenance and w is the number of weeks. c = 340 + 28w c = 280 + 32w The solution to the system is (15, 760). Which interpretation correctly describes the solution to the system of equations? A. At 15 wk of pool cleanings, a customer will reach the maximum payment of $535. B. At 15 wk, the cost for pool maintenance is the same for both companies. Both companies will charge $760. C. Paul’s Pool Service charges more money in the fifteenth week by charging $760. D. Nelson’s Pool Service charges more money in the fifteenth week by charging $760.

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

@phi @iGreen @waterineyes @WWWHHHAAATTT?

OpenStudy (wwwhhhaaattt?):

Im sorry but I cant help. Im more of game/music person so I cant help. SORRY.

OpenStudy (igreen):

I answered this question today..xD

OpenStudy (igreen):

(15, 760) means that in 15 weeks, both companies will charge $760.

OpenStudy (igreen):

@NunChuckBoy3101

OpenStudy (igreen):

So which option do you think it is? @NunChuckBoy3101

OpenStudy (anonymous):

i tagged you in a diiferent one please go there

OpenStudy (anonymous):

ok i think its b

OpenStudy (anonymous):

@iGreen